OpenStudy (mayriperez5):

If m∠4 = (3x)° and m∠8 = (x + 40)°, what is the measure of ∠4? ∠4 and ∠8 are vertical.

OpenStudy (aabomosalam1998):

Use the fact that vertical angles are congruent. m∠4=m∠8=(3x)°=(x+40)° 2x°=40°, x=20° Hence m∠4=(3*20)°=60°

OpenStudy (danjs):

see if you have 2 lines crossing , the intersection has 4 angles, and the opposite ones are the same measure